Find a numerical example of cost-volume-profit (CVP) analysis, and analyze how CVP analysis is used for decision making?
Answer #1

CVP analysis assists management in knowing the relationship between cost and revenue to generate profit. With cost-volume-profit (CVP) analysis, the management can better understand the overall performance and determine what units it should sell to break even or to attain a certain level of profit. The breakeven point is the point at which sales volume equal total expenses (both fixed and variable). Therefore CVP analysis helps decision makers to understand the effect of a change in price, sales volume, and variable cost on the profit of an entity while taking fixed cost as constant. The computing of Cost volume profit analysis formula is as below:

No. of units sold * Price per unit = No. of units sold * Variable cost per unit + Fixed cost + Profit

For example: If management wants to make an annual profit of $100,000 from the sale of appliances. The production capacity is 10,000 units; fixed cost is $30,000 and variable cost per unit is $30.

Based on the above information let’s plug the numbers in CVP formula; we get:

10000*price per unit = (10,000 * 30) + $30,000 + $100,000

Price per unit= $43

Thus, with CVP analysis the management will know that it have to price its product $43, thus CVP can help the management with decision making.

Furthermore, if the managers want to know the breakeven point it can be computed as:

Break Even Point= (Fixed Cost/Variable Cost per unit) = 30,000 / 30 = 1,000 units

Thus at 1,000 units the company's sales volume equal total expenses (both fixed and variable).